OpenStudy (anonymous):

Write the equation of the ellipse in standard form 36 x + 4 y = 144

OpenStudy (hoblos):

but 36 x + 4 y = 144 is an equation of a straight line!!

OpenStudy (anonymous):

@hoblos but for ellipse the eq is \[(x/a ^{2})+(y/b ^{2})=1\]

OpenStudy (hoblos):

yes i know... but what is the ellipse you need the equation of??

OpenStudy (anonymous):

(x/2²)+(y/6²)=1 Is this what you mean?